OpenStudy (anonymous):

∫(dx)/[x/(x^2-1)^2],solve using partial fraction

OpenStudy (anonymous):

can jus tell me the final ans.

OpenStudy (anonymous):

jimmy.r u solving the qn?

OpenStudy (anonymous):

i'm trying - i'm a bit rusty at these - i,m consulting the text book!!

OpenStudy (anonymous):

so might be a while

OpenStudy (anonymous):

ok.I'll wait

OpenStudy (anonymous):

jimmy, r u working out the solution?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

i've got 4 fractions so 4 unknowns - the algebra is horrendous - there must be an easier way to do this - ill try wolframappha

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

but what is ur final ans?

OpenStudy (anonymous):

i've split it up into four fractions so 4 unknowns - very messy - there must be an easier way - i'm checking it out on wolfram alpha

OpenStudy (anonymous):

my messages arent getting thru

OpenStudy (anonymous):

yup,wat jimmy?

OpenStudy (anonymous):

ok. anyone with final ans?

OpenStudy (anonymous):

i'm trying wolframalpha but its coming up with error messages all the time

OpenStudy (anonymous):

ok

OpenStudy (zarkon):

is this your integral? ∫(dx)/[x/(x^2-1)^2] \[\int\frac{dx}{\frac{x}{(x^2-1)^2}}\]? (Why doesn't thins thing like the negative sign??)

OpenStudy (anonymous):

yes,right

OpenStudy (anonymous):

math - i gotta go - sorry i couldn't help

OpenStudy (anonymous):

ok

OpenStudy (zarkon):

so it can be written \[\int\frac{(x^2-1)^2}{x}dx\]

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

numerator is on e denominator, that means 1 is on e numerator

OpenStudy (anonymous):

denominator has[x[x^{2}-1]^{2}\]

OpenStudy (zarkon):

\[\int\frac{1}{x(x^2-1)^2}dx\]

OpenStudy (anonymous):

yes ur right!

OpenStudy (zarkon):

ok..give me a min and I'll give you an answer

OpenStudy (anonymous):

ok.

OpenStudy (zarkon):

this is from my ti-nspire \[\frac{-\left[(x^2-1)\ln|x^2-1|-2(x^2-1)\ln|x|+1\right]}{2(x^2-1)}\]

OpenStudy (zarkon):

or from wolfram http://www.wolframalpha.com/input/?i=integral [1%2F%28x*%28x^2-1%29^2%29]+&t=ietb01

OpenStudy (anonymous):

is this ur ans?

OpenStudy (zarkon):

http://alturl.com/vfbsr

OpenStudy (zarkon):

my nspire combines everything..the wolfram answer is nicer looking

OpenStudy (anonymous):

ok. thanks